# LlaSMol: Advancing Large Language Models for Chemistry with a Large-Scale, Comprehensive, High-Quality Instruction Tuning Dataset

### v1

@@ Abstract
Chemistry plays a crucial role in many domains, such as drug discovery and material science. While large language models (LLMs) such as GPT-4 exhibit remarkable capabilities on natural language processing tasks, existing work shows their performance on chemistry tasks is discouragingly low. In this paper, however, we demonstrate that our developed LLMs can achieve very strong results on a comprehensive set of chemistry tasks, outperforming the most advanced GPT-4 across all the tasks by a substantial margin and approaching the SoTA task-specific models. The key to our success is a large-scale, comprehensive, high-quality dataset for instruction tuning named SMolInstruct. It contains 14 meticulously selected chemistry tasks and over three million high-quality samples, laying a solid foundation for training and evaluating LLMs for chemistry. Based on SMolInstruct, we fine-tune a set of open-source LLMs, among which, we find that Mistral serves as the best base model for chemistry tasks. We further conduct analysis on the impact of trainable parameters, providing insights for future research.111Dataset and model will be released at https://github.com/OSU-NLP-Group/LLM4Chem
The Ohio State University
{yu.3737, baker.3239, chen.8484, ning.104, sun.397}@osu.edu
https://osu-nlp-group.github.io/LLM4Chem/
@@ 1 Introduction
On the other hand, large language models (LLMs) such as GPT-4 (Achiam et al., 2023), Llama series (Touvron et al., 2023a, b), and Mistral (Jiang et al., 2023) have emerged as general-purpose foundation models and demonstrate remarkable abilities on various natural language processing tasks (Chang et al., 2024; Thirunavukarasu et al., 2023; Yue et al., 2023; Zhang et al., 2023; Deng et al., 2023). However, when applied to chemistry tasks, LLMs show only limited capabilities (Jablonka et al., 2022; Guo et al., 2023; Hatakeyama-Sato et al., 2023). For example, Guo et al. (2023) conducted evaluations on eight chemistry tasks and observed that while GPT-4 outperforms other closed- and open-source LLMs, its performance is not comparable with that of task-specific deep learning models. Particularly, they found that GPT models perform poorly when a precise understanding of SMILES (Weininger, 1988), a widely used textual representation for molecules, is required. In addition to directly applying pretrained LLMs, Fang et al. (2023) fine-tuned LLMs on an instruction tuning dataset, but their performance remains very low, far behind the state-of-the-art (SoTA) models designed and trained for specific tasks.
@@
Given the discouraging results thus far, some critical questions arise: Are LLMs actually able to effectively perform chemistry tasks? Or, Are they fundamentally limited for chemistry? In this paper, we demonstrate that our developed LLMs can achieve very strong results on a comprehensive set of chemistry tasks, outperforming the most advanced GPT-4 across all the tasks by a substantial margin and approaching the SoTA task-specific models.
@@
What makes such LLMs possible? First, we carefully construct a large-scale, comprehensive, and high-quality dataset for instruction tuning named SMolInstruct. We meticulously incorporate tasks with meaningful applications, collect data from diverse data sources, and apply rigorous scrutiny for quality control. The resulting dataset consists of 14 tasks and over 3M samples, laying a solid foundation for training and evaluating LLMs for chemistry tasks. The tasks and examples in the dataset are illustrated in Figure 1. Furthermore, we build a series of LLMs for chemistry named LlaSMol by fine-tuning four open-source LLMs namely Galactica, Llama 2, Code Llama, and Mistral, on SMolInstruct with LoRA (Hu et al., 2022).
We conduct a comprehensive comparison between LlaSMol and various LLMs, which reveals that LlaSMol outperforms existing LLMs including GPT-4 across all the tasks by a substantial margin. Interestingly, we found that the Mistral-based model can significantly lead the performance among all the LlaSMol models, showcasing Mistral’s great potential on chemistry. Moreover, by tuning only a small fraction of parameters once for all the tasks, LlaSMol can achieve comparable results to SoTA task-specific models that are designed and trained specifically for each individual task. Our further exploration reveals that adding trainable parameters could lead to huge performance boost. This suggests that larger scale training combined with more sophisticated base models could promisingly allow LLMs to match or surpass SoTA task-specific models on chemistry.
In answering the critical questions raised earlier, our findings underscore the great potential of LLMs to effectively perform chemistry tasks. While there is still room for further improvement, our LlaSMol series can serve as a strong set of foundation models for chemistry in the future.
@@ 2 Related Work
Task-specific Models for Chemistry. In recent years, many deep learning models have been developed to tackle different chemistry tasks. For example, Molecular Transformer (Schwaller et al., 2019) and RSMILES (Zhong et al., 2022) formulate forward synthesis and retrosynthesis prediction as sequence-to-sequence translation problems. Chemformer (Irwin et al., 2022) pretrains a transformer model on a large-scale SMILES dataset and fine-tunes it for various downstream tasks, such as forward synthesis and property prediction. Uni-Mol (Zhou et al., 2023) incorporates 3D information of molecules into the pretraining of a transformer model and fine-tunes it for downstream tasks. MolT5 (Edwards et al., 2022) first pretrains a T5 model on both SMILES and natural language, and then fine-tunes it to translate SMILES into natural language (i.e., molecule captioning) or vice versa (i.e., molecule generation). Despite their effectiveness, these models operate on single tasks and therefore cannot harness knowledge shared across diverse chemistry tasks like LLMs.
@@
LLMs for Chemistry. Recent efforts have integrated LLMs with chemistry to solve key chemistry problems, such as molecule property prediction and retrosynthesis. These efforts can be divided into two categories: (1) benchmark studies, (2) fine-tuning LLMs with new datasets. Multiple benchmark studies have evaluated (White et al., 2023; Guo et al., 2023; Jablonka et al., 2023; Liu et al., 2023) the capabilities and limitations of different off-the-shelf LLMs, such as GPT-4 and Llama, on chemistry problems. For example, Guo et al. (2023) finds that these LLMs do not perform well on chemistry tasks and often produce chemically implausible outputs. These findings highlight the need for further efforts to improve LLMs via fine-tuning for chemistry tasks.
@@
To improve LLMs for chemistry tasks, multiple instruction tuning datasets specific to chemistry have been developed. Mol-Instructions (Fang et al., 2023) consists of 1.3M instructions for multiple small molecule tasks. However, according to our results (Section 5.3), fine-tuning on their dataset does not consistently improve LLMs’ performance when compared to LLMs without fine-tuning. Drugchat (Liang et al., 2023) collects an instruction tuning dataset with 10.8K drug molecules along with 143K instructions regarding their drug-specific properties. MolOpt-Instructions (Ye et al., 2023) consists of instructions with 1M molecule pairs for molecule optimization on six properties, in which each pair has similar molecules with different properties. Compared with these datasets, SMolInstruct is much larger and covers a more diverse and comprehensive set of chemistry tasks. This could enable LLMs to better understand molecule representations and learn chemistry knowledge across tasks.
3 Preliminaries

### v4

@@ Abstract
Chemistry plays a crucial role in many domains, such as drug discovery and material science. While large language models (LLMs) such as GPT-4 exhibit remarkable capabilities on natural language processing tasks, existing research indicates that their performance on chemistry tasks is discouragingly low. In this paper, however, we demonstrate that our developed LLMs can achieve very strong results on a comprehensive set of chemistry tasks, outperforming the most advanced GPT-4 and Claude 3 Opus by a substantial margin. To accomplish this, we propose SMolInstruct, a large-scale, comprehensive, and high-quality dataset for instruction tuning. It contains 14 selected chemistry tasks and over three million samples, laying a solid foundation for training and evaluating LLMs for chemistry. Using SMolInstruct, we fine-tune a set of open-source LLMs named as LlaSMol, among which, we find that Mistral serves as the best base model for chemistry tasks. Our analysis further demonstrates the critical role of the proposed dataset in driving the performance improvements.111Our dataset and models can be found at https://osu-nlp-group.github.io/LLM4Chem/.
@@ 1 Introduction
On the other hand, large language models (LLMs) such as GPT-4 (OpenAI, 2023), Llama series (Touvron et al., 2023a; b), and Mistral (Jiang et al., 2023) have emerged as general-purpose foundation models and demonstrate remarkable abilities on various natural language processing tasks (Chang et al., 2024; Thirunavukarasu et al., 2023; Yue et al., 2023; Zhang et al., 2023; Deng et al., 2023). However, when applied to chemistry tasks, LLMs show only limited capabilities (Jablonka et al., 2022; Guo et al., 2023; Hatakeyama-Sato et al., 2023). For example, Guo et al. (2023) conducted evaluations on eight chemistry tasks and observed that while GPT-4 outperforms other closed- and open-source LLMs, its performance is far from that of task-specific deep learning models. Particularly, they found that GPT models perform poorly when a precise understanding of SMILES (Weininger, 1988), a widely used textual representation for molecules, is required. In addition to directly applying pretrained LLMs, Fang et al. (2023) fine-tuned LLMs on an instruction tuning dataset, but their performance remains very low, far behind the state-of-the-art (SoTA) models designed and trained for specific tasks.
@@
Given these discouraging results, some critical questions arise: Are LLMs actually able to effectively perform chemistry tasks? Or, Are they fundamentally limited for chemistry? In this paper, we demonstrate that our developed LLMs can achieve very strong results on a comprehensive set of chemistry tasks, substantially outperforming the most advanced GPT-4 OpenAI (2023) and Claude 3 Opus Anthropic (2024).
@@
What makes such LLMs possible? First, we construct a large-scale, comprehensive, and high-quality dataset for instruction tuning named SMolInstruct. We incorporate tasks with meaningful applications, collect data from diverse data sources, and apply rigorous scrutiny for quality control. The resulting dataset consists of 14 tasks (illustrated in Figure 1) and over 3M samples, laying a solid foundation for training and evaluating LLMs for chemistry tasks. Based on the dataset, we build a series of LLMs for chemistry named LlaSMol by fine-tuning four open-source LLMs namely Galactica, Llama 2, Code Llama, and Mistral, on SMolInstruct with LoRA (Hu et al., 2022).
We conduct comprehensive experiments to evaluate our models and explore their insights, yielding some interesting findings. Firstly, among the four LlaSMol models, the Mistral-based model surpasses others by a substantial margin, showcasing the considerable influence of base models on downstream chemistry tasks. Moreover, contrast to claims made in previous work (Fang et al., 2023), using SMILES as the molecular representation achieves sufficient validity of generated molecules and better performance compared to using SELFIES (Krenn et al., 2019). Furthermore, employing canonicalized SMILES during model training and applications can alleviate learning burdens and increase performance. Finally, while instruction tuning can inject chemistry task-related knowledge into models, the dataset plays a crucial role. Our experiments demonstrate that training on our SMolInstruct leads to substantially better performance compared to training on previous dataset, emphasizing the contribution of the proposed dataset. Although LlaSMol models do not yet surpass state-of-the-art (SoTA) task-specific models that are designed and trained specifically for each individual task, they approach SoTA performance with only 0.58% of parameters being fine-tuned, suggesting their great potential for further improvements and to serve as strong foundation models for the field.
@@ 2 Related Work
Task-specific Models for Chemistry. In recent years, many deep learning models have been developed to tackle different chemistry tasks. For example, Molecular Transformer Schwaller et al. (2019) and RSMILES Zhong et al. (2022) formulate forward synthesis and retrosynthesis prediction as sequence-to-sequence translation problems. Chemformer Irwin et al. (2022) pretrains a transformer model on a large-scale SMILES dataset and fine-tunes it for various downstream tasks, such as forward synthesis and property prediction. MolT5 Edwards et al. (2022) first pretrains a T5 model on both SMILES and natural language, and then fine-tunes it to translate SMILES into natural language (i.e., molecule captioning) or vice versa (i.e., molecule generation). Graph neural networks (GNNs), which directly leverage the graph structure of the molecule Wang et al. (2023), have also shown promise in many chemistry applications, such as property prediction Yang et al. (2019); Han et al. (2023), retrosynthesis Chen et al. (2023); Somnath et al. (2021), and molecule optimization Chen et al. (2021); Zhang et al. (2022b). Recent studies Zhou et al. (2023); Zhang et al. (2022a) have shown the promise of leveraging equivariant representations of molecular 3D structures for chemistry tasks, such as property prediction Zhou et al. (2023) and docking Zhang et al. (2022a). Uni-Mol Zhou et al. (2023) incorporates this 3D information into the pretraining of a transformer model and fine-tunes it for downstream tasks. Despite their effectiveness, these models operate on single tasks and therefore cannot harness knowledge shared across diverse chemistry tasks like LLMs.
@@
LLMs for Chemistry. Recent efforts have integrated LLMs with chemistry to solve key chemistry problems, which can be divided into two categories: (1) benchmark studies, and (2) fine-tuning LLMs with new datasets. Multiple benchmark studies White et al. (2023); Guo et al. (2023); Jablonka et al. (2023); Liu et al. (2023a) have evaluated the capabilities and limitations of different off-the-shelf LLMs, such as GPT-4 and Llama, on chemistry problems. For example, Guo et al. (2023) finds that these LLMs do not perform well on chemistry tasks and often produce chemically implausible outputs. These findings highlight the need for further efforts to improve LLMs via fine-tuning for chemistry tasks.
@@
To improve LLMs for chemistry, multiple instruction tuning datasets have been developed. Mol-Instructions Fang et al. (2023) consists of 1.3M instructions for multiple small molecule tasks. However, fine-tuning on the dataset does not significantly improve LLMs’ performance (Section 4.3). Drugchat Liang et al. (2023) collects an instruction tuning dataset on drug properties with 10.8K drug molecules. MolOpt-Instructions Ye et al. (2023) consists of instructions with 1M molecule pairs for molecule optimization on six properties, in which each pair has similar molecules with different properties. Recent works also develop 2D or 3D molecular graph-centric datasets and integrate the graph understanding ability into LLMs Liu et al. (2023b); Cao et al. (2023); Li et al. (2024). Compared with these datasets, SMolInstruct is much larger and covers a more diverse and comprehensive set of chemistry tasks, which enables LLMs to better understand molecule representations and learn chemistry knowledge across tasks.
3 SMolInstruct
This section introduces our proposed dataset SMolInstruct and its construction. Readers may refer to Appendix A for preliminaries and background.
3.1 Overview of SMolInstruct
SMolInstruct is a large-scale instruction tuning dataset that centers around small molecules. It contains 14 chemistry tasks, illustrated in Figure 1.
(1) We include four name conversion tasks, namely converting IUPAC name to molecular formula (NC-I2F), converting IUPAC name to SMILES (NC-I2S), converting SMILES to molecular formula (NC-S2F), and converting SMILES to IUPAC name (NC-S2I). They are designed to enable deep understanding of molecular structures and representations, which should serve as the fundamental knowledge for chemistry LLMs.
(2) Additionally, six property prediction tasks (Wu et al., 2018) are integrated, including PP-ESOL for water solubility (Mobley & Guthrie, 2014), PP-Lipo for octanol/water distribution coefficient (Poole & Poole, 2003), PP-BBBP for blood-brain barrier penetration (Martins et al., 2012), PP-ClinTox for toxicity to human body (Gayvert et al., 2016), PP-HIV for HIV replication inhibition (Institute, 2004), and PP-SIDER for side effects of drugs (Kuhn et al., 2015). These involved properties are crucial especially for drug development.
SMolInstruct contains 3.3M samples. Each sample is a query-response pair, where the query describes a task and any task-specific information (e.g., input molecule, textual description, etc.), and the response is a sentence containing the answer to the queried task. For all the tasks, unless explicitly defined in the tasks (NC-I2F, NC-I2S, NC-S2F, and NC-S2I), we use SMILES as the default representation for molecules, but also provide the SELFIES (Krenn et al., 2019) representation.
3.2 SMolInstruct Construction
We construct the SMolInstruct dataset by following a four-step pipeline: data collection, quality control, data splitting, and instruction construction.
Data Collection. After consulting domain experts and pinpointing the set of meaningful tasks (summarized in Section 3.1), we collect data for these tasks from various sources, as listed in Table 5. Specifically, for the name conversion tasks (NC-I2F, NC-I2S, NC-S2F, and NC-S2I), we leverage PubChem222https://pubchem.ncbi.nlm.nih.gov/ (Kim et al., 2019), one of the most comprehensive molecule databases. Within this database, we randomly select a large set of molecule entries, and extract their IUPAC names, SMILES representations, and molecular formulas. This obtained data is then re-organized as input-output pairs for the tasks. For molecular description-related tasks (MC and MG), we utilize a combination of ChEBI-20 (Edwards et al., 2021; 2022) and Mol-Instructions (Fang et al., 2023), as they both contain high-quality molecule-text paired data. For property prediction tasks (PP-ESOL, PP-Lipo, PP-BBBP, PP-ClinTox, PP-HIV, and PP-SIDER), we employ the well-established MoleculeNet datasets (Wu et al., 2018). We select the 6 datasets from MoleculeNet that represent the essential properties for real-world applications such as drug discovery. For chemical reaction tasks (FS and RS), we collect the reaction data from USPTO-full (Lowe, 2017), which is an extensive collection encompassing over 1M reaction samples extracted from U.S. patents. All the aforementioned datasets are also widely used in previous studies (He et al., 2021; Zhong et al., 2022; Edwards et al., 2022; Irwin et al., 2022; Chen et al., 2023; Zhou et al., 2023).
Quality Control. To guarantee high quality, we apply rigorous scrutiny. The collected data contains many problematic and low-quality samples, which can be roughly categorized into the following three types, along with our curation methods: (1) Chemically invalid SMILES. Numerous SMILES strings are chemically invalid (e.g., deviating from the SMILES grammar, or violating chemical valence). To address this issue, we employ RDKit (RDKit, 2023), a widely used toolkit for cheminformatics, to parse molecules and detect errors. (2) Wrong or inaccurate information. Based on manual check, we observed wrong and inaccurate information recorded in the data. For instance, within the USPTO-full dataset (Lowe, 2017), we identify and correct mislabeled reactants and reagents in chemical reactions by comparing their atom mappings with products. For the MC and MG tasks, we filter out those textual descriptions that lack pertinent, molecule-specific information, with a set of rules based on wording patterns, lengths and keywords. For PP-SIDER, we eliminate disorders with ambiguous names that could impede the creation of precise and comprehensible instructions. (3) Duplicated samples. We detect and remove them.
Data Splitting. Data splitting for multi-task datasets requires careful handling in order to avoid data leakage across tasks. For instance, FS and RS are a pair of reverse tasks, so data leakage occurs when the training set contains an FS sample for a certain chemical reaction and the test set has an RS sample for the same reaction. This can lead to biased evaluation. Therefore, we identify sample pairs across related tasks (FS and RS, MC and MG, and the four NC tasks) that correspond to the same molecules/reactions, and ensure that matched samples are placed together in either training or evaluation set. Moreover, some samples may share the same input but have different outputs. For instance, in the RS task, one product (the same input) may be synthesized from multiple sets of reactants (different outputs). If these samples are placed into both training and test set, it may lead to exaggerated performance. Therefore we ensure that samples with identical inputs are placed together either in or outside of the test set. Additionally, to achieve fair comparisons with Mol-instructions (Fang et al., 2023), for tasks shared between the two datasets (MC, MG, FS, and RS), we ensure that their training examples are not included in the test set of SMolInstruct, allowing for a direct evaluation of their models on our test set. Following these necessary limitations, samples are randomly split into training/validation/test set, except for PP task samples that undergo a scaffold splitting following the canonical method (Wu et al., 2018).
For more details of dataset construction, please refer to Section B.2.
3.3 Merits of SMolInstruct
(1) Large-Scale. SMolInstruct consists of 3.3M samples and 1.6M distinct molecules, with a diverse range of sizes, structures, and properties (see Section B.1), showcasing an extensive coverage of diverse chemical knowledge.
(2) Comprehensive. SMolInstruct contains 4 types of chemical tasks (14 tasks in total), emerging as the most comprehensive instruction tuning dataset for small molecules. Notably, the tasks are meticulously selected to build a strong chemistry foundation model and to adapt to real-world applications.
(3) High-Quality. Rigorous processing steps have been implemented to exclude problematic and low-quality samples. Along with careful data splitting and canonicalization of SMILES representations, SMolInstruct stands as a high-quality resource valuable for future research.
A detailed introduction and statistics of the SMolInstruct dataset can be found in Appendix B. For a comparison with the previous work, Mol-Instructions (Fang et al., 2023), please refer to Appendix C.
4 Experiments
4.1 Our LlaSMol Models
By fine-tuning base models on the proposed SMolInstruct dataset, we create LLMs capable of performing chemistry tasks, which we name LlaSMol (Large language models on Small Molecules). Specifically, we extensively consider four different LLMs as our base models, namely Galactica 6.7B (Taylor et al., 2022), Llama 2 (Touvron et al., 2023b) 7B, Code Llama (Roziere et al., 2023) 7B, and Mistral (Jiang et al., 2023) 7B, where Galactica is trained for scientific applications and has already been exposed to chemistry-related data during its pretraining, Llama 2 and Mistral are general-purpose LLMs, while Code Llama is based on Llama 2 and trained for code. We conduct instruction tuning on the proposed SMolInstruct dataset, and name the resulting models as LlaSMol, LlaSMol, LlaSMol, and LlaSMol, respectively. All the LlaSMol models are trained with LoRA (Hu et al., 2022), which is applied to all weight matrices in the self-attention and feedforward neural network (FFN) modules with lora_r and lora_alpha set to 16. The fine-tuning process utilizes the Huggingface Transformers library (Wolf et al., 2020). Training spans three epochs, employing the 8-bit AdamW optimizer, a learning rate of 1e-4, and a cosine scheduler. The input length for training is set to 512, which covers 99.7% of the samples. During inference, we adopt beam search as the generation strategy for simplicity.
4.2 Experimental Setup
(1) LLMs without fine-tuning on SMolInstruct. This type includes our four base models, namely Galactica (Taylor et al., 2022), Llama 2 (Touvron et al., 2023b), Code Llama (Roziere et al., 2023), Mistral (Jiang et al., 2023). we also benchmark against GPT-4 (OpenAI, 2023) and the more recent Claude 3 Opus (Anthropic, 2024), the current state-of-the-art (SoTA) LLMs333Due to resource limitations, we evaluate GPT-4 and Claude 3 Opus on at most 500 test samples for each task.. For Llama 2, Code Llama, and Mistral, we use 1-shot, due to their poor instruction following ability; for GPT-4, we report its results under a zero-shot setting, as GPT-4 performs best on this setting in our experiments (Appendix E); for Claude 3 Opus, we report its zero-shot results as well. We also include two LLMs tuned specifically for chemistry tasks: Molinst, a Llama 2 model tuned on the Mol-Instructions dataset by Fang et al. (2023), which shares the training tasks of MC, MG, FS, and RS with LlaSMol; and ChemLLM (Zhang et al., 2024), an LLM for chemistry proposed concurrently to our work.
(2) SoTA task-specific models. To provide a comprehensive view of LlaSMol’s performance, we present results from SoTA task-specific models. For NC-I2S and NC-S2I, we compare with STOUT (Rajan et al., 2021), an encoder-decoder model trained on SMILES-IUPAC name paired data. For NC-S2F, a task achievable with a fixed algorithm, we implement a program with RDKit (RDKit, 2023), a widely used Python toolkit for cheminformatics, and report its results. For NC-I2F where no dedicated models exist, we construct a baseline called STOUT+RDKit by aggregating STOUT for I2S conversion and RDKit for S2F conversion. For the PP tasks, our compared model is Uni-Mol (Zhou et al., 2023). It incorporates molecular 3D representations and follows a pretraining and fine-tuning paradigm. Following its original settings, we fine-tune the model on our SMolInstruct dataset with its pretrained checkpoint. In the case of MC and MG, we compare with MolT5 (Edwards et al., 2022) and directly use their released checkpoint. The reasons why we do not use our re-trained model are: (1) we were unable to reproduce results close to those reported in the paper as no original code was provided; and (2) we take great care to ensure that our test set is devoid of training examples used by MolT5, ensuring fairness in the evaluation. Lastly, regarding FS and RS, we re-train RSMILES (Zhong et al., 2022) and Molecular Transformer (Schwaller et al., 2019) for the two tasks, respectively, following their reported settings. Both of the models are transformer encoder-decoder models (Vaswani et al., 2017), specifically adapted for the FS and RS tasks.
Evaluation Metrics. We employ metrics commonly used in previous work (Schwaller et al., 2019; Zhong et al., 2022; Fang et al., 2023; Zhou et al., 2023; Chen et al., 2023), which include: (1) Exact Match (EM), indicating the proportion of predicted results that exactly match the gold standards. (2) Fingerprint Tanimoto Similarity (FTS), quantifying structural similarities between molecules using Tanimoto similarities of their Morgan fingerprints (Morgan, 1965). (3) METEOR score, a comprehensive text-based metric considering both exact matches and semantic similarity (Lavie & Agarwal, 2007) for the MC task. (4) Root Mean Square Error (RMSE), measuring the square root of the average squared differences between predicted and actual values for the PP-ESOL and PP-Lipo tasks (5) Accuracy (Acc), the ratio of correct predictions for the binary classification tasks (PP-BBBP, PP-ClinTox, PP-HIV, and PP-SIDER). (6) Validity (Valid), the ratio of valid predictions following SMILES grammar and chemical valence rules for tasks with SMILES outputs (NC-I2S, MG, FS, and RS). For all the metrics except RMSE, higher values indicate better performance.
4.3 Main Results
[b]
| Model | NC | PP | |||||||||
| I2F | I2S | S2F | S2I | ESOL | Lipo | BBBP | Clintox | HIV | SIDER | ||
| EM | EM | Valid | EM | EM | RMSE | RMSE | Acc | Acc | Acc | Acc | |
| Task-Specific, Non-LLM Based Models | |||||||||||
| SoTA | |||||||||||
| Existing LLMs without fine-tuning on SMolInstruct | |||||||||||
| GPT-4 | |||||||||||
| Claude 3 Opus | |||||||||||
| Galactica | |||||||||||
| Llama 2 | |||||||||||
| Code Llama | |||||||||||
| Mistral | |||||||||||
| Molinst (chemistry LLM) | |||||||||||
| ChemLLM (chemistry LLM) | |||||||||||
| Our LlaSMol Series | |||||||||||
| LlaSMol | |||||||||||
| LlaSMol | |||||||||||
| LlaSMol | |||||||||||
| LlaSMol | |||||||||||
(1) Among all the LLMs, our LlaSMol models demonstrate the best performance, underscoring the effectiveness of the proposed SMolInstruct dataset and fine-tuning. Specifically, compared to the base models (Galactica, Llama 2, Code Llama, and Mistral), LlaSMol models exhibit substantial performance improvements, which highlights the effectiveness of SMolInstruct in enhancing the understanding of molecular representations and the task-related knowledge, and signifies the effective learning of chemistry-related tasks by LLMs. Furthermore, LlaSMol substantially outperforms GPT-4 on all the tasks and Claude 3 Opus on most tasks, despite their larger parameter size. LlaSMol also surpasses the two chemistry LLMs namely ChemLLM444Since its dataset and evaluation details are not available, we cannot provide more analysis., which is similarly trained on chemistry instruction data. and Molinst. Notably, LlaSMol, which uses the same base model and LoRA setting as Molinst, outperforms it even on the shared training tasks (MC, MG, FS, and RS). This finding highlights the benefits of our dataset.
(2) Our four LlaSMol models show substantial differences in their performance, emphasizing the considerable impact of base models on downstream tasks. Despite sharing identical training, inference settings, and comparable model sizes, LlaSMol consistently outperforms LlaSMol by a substantial margin, highlighting Mistral’s potential on chemistry tasks. In addition, LlaSMol exhibits better performance than LlaSMol on most tasks, indicating a potential synergy between programming language knowledge in Code Llama and molecular representations. Furthermore, LlaSMol outperforms LlaSMol, and LlaSMol in most cases, suggesting the benefits of pretraining on chemistry-related documents.
(3) Although LlaSMol models do not outperform SoTA models, they demonstrate considerable potential for further improvements. Specifically, LlaSMol surpasses the SoTA models on PP-Clintox and PP-SIDER, but has yet to achieve the success on other tasks. However, LlaSMol has greatly narrowed the performance gap between LLMs and SoTA task-specific models, compared to previous efforts (Fang et al., 2023; Zhang et al., 2024). Remarkably, LlaSMol attains such performance with only a small proportion of its parameters fine-tuned (approximately 41.9M, 0.58% of its parameters). As shown in Section F.2, increasing the number of trainable parameters can substantially boost performance, suggesting that LlaSMol has immense potential to surpass task-specific models through more extensive fine-tuning and serve as a strong foundation model for chemistry applications.
| Model | MC | MG | FS | RS | ||||||
| METEOR | EM | FTS | Valid | EM | FTS | Valid | EM | FTS | Valid | |
| Task-Specific, Non-LLM Based Models | ||||||||||
| SoTA | ||||||||||
| Existing LLMs Without Fine-Tuning on SMolInstruct | ||||||||||
| GPT-4 | ||||||||||
| Claude 3 Opus | ||||||||||
| Galactica | ||||||||||
| Llama 2 | ||||||||||
| Code Llama | ||||||||||
| Mistral | ||||||||||
| Molinst (chemistry LLM) | ||||||||||
| ChemLLM (chemistry LLM) | ||||||||||
| Our LlaSMol Series | ||||||||||
| LlaSMol | ||||||||||
| LlaSMol | ||||||||||
| LlaSMol | ||||||||||
| LlaSMol | ||||||||||
4.4 Ablation Study
To investigate the advantages of SMolInstruct, we conduct an ablation study by comparing LlaSMol with the following variants: (1) w/o canonical, which uses uncanonicalized SMILES, to examine the benefits of canonicalization. (2) using SELFIES, which uses SELFIES Krenn et al. (2019) instead of SMILES to explore their differences. (3) train on Mol-Instructions, which is trained on Mol-Instructions (Fang et al., 2023), to compare the performance improvements of our dataset against the previously proposed dataset.
[b]
| Model | NC | PP | |||||||||
| I2F | I2S | S2F | S2I | ESOL | Lipo | BBBP | Clintox | HIV | SIDER | ||
| EM | EM | Valid | EM | EM | RMSE | RMSE | Acc | Acc | Acc | Acc | |
| LlaSMol | |||||||||||
| w/o canonical | 88.5 | 67.2 | 93.4 | 24.5 | 1.224 | 1.072 | 71.6 | 96.8 | 70.3 | ||
| using SELFIES | 86.9 | 47.7 | 100.0 | 94.7 | 19.7 | 1.456 | 1.106 | 69.5 | 91.7 | 96.5 | 64.4 |
| train on Mol-Instructions | 0.0 | 0.0 | 75.2 | 0.0 | 0.0 | 4.416 | 2.282 | 0.0 | 0.0 | 2.6 | 0.4 |
| Model | MC | MG | FS | RS | ||||||
| METEOR | EM | FTS | Valid | EM | FTS | Valid | EM | FTS | Valid | |
| LlaSMol | ||||||||||
| w/o canonical | 0.457 | 16.8 | 60.2 | 99.1 | 53.7 | 80.8 | 99.9 | 23.8 | 67.4 | 99.9 |
| using SELFIES | 0.466 | 16.2 | 58.6 | 99.9 | 40.4 | 74.0 | 100.0 | 25.6 | 66.0 | 99.9 |
| train on Mol-Instructions | 0.195 | 6.1 | 46.1 | 88.2 | 3.9 | 37.1 | 78.3 | 7.4 | 52.6 | 76.7 |
The results in Table 3 and Table 4 lead to the following observations: (1) The “w/o canonical” model underperforms LlaSMol on most tasks, with a substantial performance drop on FS and RS. This suggests that canonicalizing SMILES can reduce learning difficulty and improve performance. As canonicalization can be easily performed using fixed algorithms before feeding into models, we recommend using canonical SMILES when training and applying LLMs for chemistry. (2) While using SELFIES slightly improves the validity of generated molecules, which aligns with the motivation behind SELFIES (Krenn et al., 2019), the validity of using SMILES is also sufficiently high. Moreover, using SELFIES results in worse performance on most tasks, possibly due to SELFIES being typically longer than SMILES, making it more difficult for the model to accurately understand and generate. Therefore, using SELFIES over SMILES may not be necessary, contrast to claims made in previous work (Krenn et al., 2019; Fang et al., 2023). (3) Despite using identical base models and training settings, the model trained on Mol-Instructions (Fang et al., 2023) performs much worse than LlaSMol trained on SMolInstruct even on the shared tasks (MC, MG, FS, and RS). This demonstrates the superiority of our dataset. A detailed comparison with Mol-Instructions can be found in Appendix C.
To gain deeper insights into the models’ performance and behavior, we conduct further analytical experiments: (1) To investigate the synergistic effects among different tasks, we evaluate models trained on a single task and models with certain tasks removed. The results demonstrate multiple-task training outperforms single-task training, indicating its benefits. However, each task generally does not heavily rely on the presence of other tasks, suggesting a degree of independence among them. (2) To investigate the influence of LoRA (Hu et al., 2022) settings, we vary the involved LoRA modules. We observe that adding LoRA modules (and trainable parameters) leads to a substantial boost in performance, indicates the models’ great potential for further improvements if with larger-scale fine-tuning. Please refer to Appendix F for more details.
5 Conclusion
While LLMs have shown promise as versatile assistants, their performance on chemistry-related tasks remains notably subpar. To address this issue, we introduces SMolInstruct, a large-scale, comprehensive, and high-quality instruction tuning dataset. It comprises 14 tasks highly relevant to real-world applications and contains over 3M rigorously curated samples. Using SMolInstruct, we develop LlaSMol, a series of LLMs for performing chemistry tasks. Our experiments demonstrate LlaSMol’s superiority over existing LLMs, and highlight SMolInstruct’s crucial role in boosting the performance. Further analytical experiments also provide significant insights towards developing LLMs for chemistry.
However, this work has the following limitations. First, the evaluations for the MC and MG tasks cannot accurately assess models’ abilities to generate chemically correct descriptions and molecules. Since the definition of molecular descriptions remain ambiguous and the available data is limited, it is challenging to assess whether the generated descriptions or molecules are accurate and correct. Second, this work does not delve into the models’ generalization capabilities beyond the trained tasks. While we recognize the importance of such capabilities, how to meaningfully test generalization abilities is nontrivial and needs careful design, which falls outside the purview of this work. Third, our models do not yet outperform SoTA task-specific models, possibly due to the small ratio of trainable parameters or suboptimal training procedures. Nevertheless, we propose a high-quality instruction tuning dataset, demonstrate its effectiveness, and gain deeper insights, which we hope can be valuable for future research. We will try to address the aforementioned limitations in our future work.
Ethics Statement
Despite our best efforts to maintain the high quality of the SMolInstruct dataset and the integrity of the LlaSMol models, we cannot guarantee that the dataset is free of inaccurate, incorrect, or harmful content, nor can we prevent the models from generating such content. Users should engage with our dataset and models at their own discretion and uphold the highest ethical standards in their use.
Acknowledgement
The authors would thank colleagues from the OSU NLP group and the OSU Ning Lab for constructive feedback. This research was supported in part by NSF IIS-2133650, NIH 1R01LM014385-01, and NSF CAREER #1942980, as well as Ohio Supercomputer Center (Ohio Supercomputer Center, 1987). The views and conclusions contained herein are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the U.S. government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notice herein.
Appendix A Preliminaries